# Synergistic Bleeding Region and Point Detection in Laparoscopic Surgical Videos

### v2

@@ Abstract
Intraoperative bleeding in laparoscopic surgery causes rapid obscuration of the operative field to hinder the surgical process and increases the risk of postoperative complications. Intelligent detection of bleeding areas can quantify the blood loss to assist decision-making, while locating bleeding points helps surgeons quickly identify the source of bleeding and achieve hemostasis in time to improve surgical success rates. In this study, we first construct a real-world laparoscopic surgical bleeding detection dataset, named SurgBlood, comprising 5,330 frames from 95 surgical video clips with bleeding region and point annotations. Accordingly, we develop a dual-task synergistic online detector called BlooDet, designed to perform simultaneous detection of bleeding regions and points in laparoscopic surgery. Our framework embraces a dual-branch bidirectional guidance design based on Segment Anything Model 2 (SAM 2). The mask branch detects bleeding regions through adaptive edge and point prompt embeddings, and the point branch leverages mask memory to induce bleeding point memory modeling and capture the direction of bleed point movement via inter-frame optical flow. By bidirectional guidance, the two branches explore potential spatial-temporal relationships while leveraging memory modeling to infer the current bleeding condition. Extensive experiments demonstrate that our baseline outperforms 12 counterparts on SurgBlood in both bleeding region and point detection. Our code and data will be publicly available.
@@ 1 Introduction
Bleeding in Laparoscopic Surgery. Minimally invasive surgery has revolutionized clinical healthcare by reducing patient trauma and accelerating postoperative recovery [36]. However, intraoperative bleeding is a common emergency that significantly impacts surgical safety and efficiency in laparoscopic surgery [13]. Rapid changes in the amount and speed of bleeding can severely obscure the surgical field, delaying the surgeon’s response and reducing the success rate of surgery. More prolonged bleeding increases the risk of organ damage and postoperative complications [11]. Therefore, utilizing computer-assisted techniques to detect bleeding regions and localize bleeding points holds significant clinical value. In one respect, detecting bleeding areas can quantify blood loss, providing timely support for intraoperative decision-making. In another respect, the precise positioning of bleeding points enables surgeons to control hemorrhage promptly to ensure surgical safety.
@@
Figure 1: (a): Illustration of bleeding detection task with samples in SurgBlood and the predictions of our solution. (b): The proposed BlooDet framework performs dual-branch bidirectional guidance for the synergistic bleeding region and point detection. Zoom in for details.
@@
Challenges for Bleeding Detection. Despite the popularity of laparoscopic surgery, automated detection of bleeding regions and bleeding points still faces numerous challenges [33]. Due to the narrow field of view under laparoscopy and unstable lighting conditions, the anatomical structures are incompletely exposed, which increases the difficulty of extracting discriminative representations. Additionally, the rapid accumulation and flow of blood can change tissue appearance and infiltrate surrounding tissues, reducing the availability of low-level visual clues and complicating the detection of bleeding regions. The bleeding points may also be buried by blood or obscured by tissues, making it difficult to quickly locate and continuously track [26]. Beyond these challenges, intelligent bleeding warning involves detecting bleeding regions and locating bleeding points during dynamic surgical procedures [15]. This requires a reliable multi-task online detector that models fine-grained spatial-temporal relationships in surgical videos for accurate predictions. Further, the lack of publicly available multi-task real bleeding datasets remains a major obstacle to progress in surgical intelligence.
@@
Proposed Benchmark. To advance research on bleeding region and point detection in surgical videos, we construct a new actual laparoscopic surgery bleeding dataset, named SurgBlood. Our dataset comprises a total of 5,330 video frames from 95 laparoscopic video clips, encompassing multiple types and intensities of bleeding during surgery. As displayed in Fig. 1(a), SurgBlood also provides pixel-level annotations of bleeding regions and bleeding point coordinates by hepatobiliary surgeons. Our dataset supports the joint detection of bleeding regions and points in video streams. We evaluate several task-relevant methods on SurgBlood to establish a comprehensive benchmark for laparoscopic intraoperative bleeding detection, driving further research in surgical assistance.
@@
Existing Methods. Various deep learning-based algorithms have been demonstrated to be effective in bleeding region detection, including applications in intracranial hemorrhage detection [18] and bleeding detection in capsule endoscopy [3]. However, most algorithms [3, 18, 9] are designed for image or keyframe analysis, lacking the ability to model temporal dependencies in surgical videos. In addition, previous methods [30, 23, 33] mainly focus on bleeding region detection, which falls short of addressing the clinical needs in locating the bleeding source. With the emergence of large vision models, Segment Anything Model 2 (SAM 2) [29] unleashed powerful visual representation capabilities for video sequence modeling. Subsequently, a series of SAM 2-based frameworks [22, 2] have been proposed for video detection in the clinical medicine domain, but have not yet been unified into a multi-task paradigm. Some multi-task frameworks [9, 23] detect both regions and keypoints concurrently, but they overlook mutual guidance for joint optimization across tasks.
@@
Our Solution. To meet the clinical demand of bleeding region and point detection, in this paper, we propose a dual-task online baseline model called BlooDet, which adopts a dual-branch bidirectional guidance structure based on video-level SAM 2 to synergistically optimize both tasks. As illustrated in Fig. 1(b), our framework consists of two branches: Mask branch and Point branch. In the mask branch, we embed an edge generator that performs multi-scale perception of spatial-temporal features with the Wavelet Laplacian filter to generate edge prompts, mitigating the problem of blurred bleeding boundaries in surgical scenes. Meanwhile, we incorporate bleeding points produced from the point branch as point prompts and combine them with edge prompts to facilitate bleeding region detection. For the point branch, considering the movement of bleeding points within the field of view is influenced by the relative motion of the camera, we leverage inter-frame optical flow and bleeding mask memory to estimate camera motion and viewpoint offsets, improving the location accuracy of the bleeding point and reducing interference caused by surrounding blood blurring. Further, we integrate mask memory features from the mask branch to enhance bleeding point location perception. By exchanging clues and co-guiding across two branches, BlooDet can exploits the underlying spatial and temporal associations between bleeding regions and points. Extensive experiments demonstrate that our approach achieves superior performance in both bleeding region and point detection tasks, e.g., 64.88% IoU for bleeding region detection and 83.69% PCK-10% for bleeding point detection.
Figure 2: Illustration of bleeding types in the SurgBlood dataset. Bleeding regions and bleeding points are labeled in yellow mask and red dot.
The main contributions of this work are four-fold: 1) We debut the intraoperative bleeding region and bleeding point detection tasks in surgical videos and contribute a real-world bleeding detection dataset, termed SurgBlood, to advance the surgical intelligent assistance community. 2) We propose a dual-task synergistic online detection model, BlooDet, for jointly detecting bleeding regions and points in surgical videos. Our framework embraces a dual-branch structure and performs co-optimization by mutual prompts and bidirectional guidance. 3) The point branch utilizes inter-frame optical flow and mask memory for point memory modeling, capturing bleeding point movement cues and providing spatial-temporal modeling. Adaptive edge and point prompting strategies are introduced in the mask branch, where the edge generator exploits multi-scale Wavelet Laplacian filter to enhance edge perception, and combined with bleeding points for mask prompt embedding. 4) We evaluate 12 task-related models and establish a comprehensive benchmark on SurgBlood. Experimental results indicate that our baseline achieves superior performance for joint bleeding region and point detection.
@@ 2 Related Work
Bleeding Region Detection in Medical Domain. Bleeding region detection has been explored across various medical scenarios, such as intracranial hemorrhage detection [18], capsule endoscopy bleeding recognition [3], and retinal hemorrhage identification [40]. Deep learning-based methods employ convolution and attention mechanisms to extract discriminative features for bleeding localization. In surgical videos, intraoperative bleeding detection presents unique challenges due to limited workspace and dynamic lighting variations. Sunakawa et al. [33] developed a semantic segmentation model for automatically recognizing bleeding regions on the anatomical structure of the liver. Nonetheless, existing methods primarily focus on mask-level bleeding detection, overlooking the localization of the bleeding source. To bridge this gap, we introduce a unified paradigm for the synergistic detection of bleeding regions and points in surgical videos.
@@
Keypoint Detection in Medical Images. Keypoint detection plays a crucial role in various clinical applications, e.g., pathological site identification and anatomical landmark localization [38, 1]. Existing keypoint detection methods are usually classified into three categories: 1) Context-aware spatial methods. This technique exploits the stability and uniformity of keypoint spatial distributions to improve localization accuracy [42, 21, 35, 27]. 2) Multi-stage learning strategies. These architectures follow a coarse-to-fine process to refine keypoint localization through gradual optimization and integrating shallow and deep layers to enhance performance [45, 8, 46]. 3) Multi-task learning frameworks, which jointly optimize medical image segmentation and keypoint detection by constructing a union network [43, 44, 12, 9]. Although current multi-task learning methods demonstrate strong performance, they neglect the mutual guidance between tasks. Therefore, we propose a collaborative dual-branch model that enhances both bleeding region and point detection via cross-task guidance.
Video Segmentation with Large Vision Models. With the advent of large vision models (LVMs) [16], video segmentation techniques have made great progress. Among these models, SAM 2 [29] has emerged as the leading framework, extending the capabilities of SAM from image segmentation to video domain. By leveraging large-scale pre-training and fine-tuning techniques, SAM 2 achieves state-of-the-art performance across a range of video tasks. LVM-based video segmentation techniques open up new possibilities in various downstream applications [28, 14, 7, 2, 47, 41]. SAM2-adapter [7] embeds adapter layers into SAM 2 to enhance the model’s flexibility, improving cross-task generalization. In the domain of surgical video analysis, SurgSAM-2 [22] utilizes SAM 2 along with a frame pruning mechanism for efficient instrument segmentation. Based on the powerful temporal modeling capabilities of SAM 2, we design a dual-task paradigm that allows for joint bleeding region and point detection in complex surgical environments.
3 Proposed SurgBlood Dataset
Figure 3: Data statistics for SurgBlood. (a) Statistical distribution of video clips; (b) Proportion of frames with bleeding region and point; (c) Distance of bleeding region and point to image center.
@@ 3 Proposed SurgBlood Dataset
Currently, there are few publicly available datasets for surgical bleeding detection. HemoSet [24] provides bleeding detection samples based on live animal robotic surgery, but it only includes 857 labeled frames for bleeding areas while ignoring bleeding point detection. To this end, we construct a brand-new dataset specifically for bleeding region and point detection in laparoscopic surgery, named SurgBlood. We provide an overview of our dataset from the following three aspects:
Data Collection. To ensure high-quality and representative data, we invited four hepatobiliary surgeons from partner hospitals to carefully select 95 video clips from 41 cholecystectomy surgical cases. Each clip covers the entire bleeding process while retaining the non-bleeding scene for approximately 3 seconds before and after the bleeding event. We collect a total of 5,330 video frames with a resolution of 1280$\times$720 from all clips using a sampling rate of 2 fps. Notably, we focus exclusively on dynamic bleeding regions within the surgical action field, as this is the critical location that directly interferes with the surgeon and contains key bleeding points. As shown in Fig. 2, there are four bleeding types by tissue location: gallbladder, cystic triangle, vessel, and gallbladder bed. The proportion of bleeding types corresponds to the frequency distribution observed in actual surgery.
@@
Data Annotation. To ensure the annotation quality of SurgBlood, the invited hepatobiliary surgeons meticulously annotate and review each video clip. During the labeling process, the surgeon uses both static frames and dynamic video sequences to label bleeding regions and bleeding point coordinates for each frame. Annotation is guided by the following principles: 1) For bleeding regions, pooled blood and inactive sparse bloodstains are not annotated. 2) For bleeding points, if the bleeding point is not obscured by tissue or instruments, its coordinates are labeled; if the bleeding point is obscured by blood, annotators trace back to the first frame where the bleeding occurred and localize the point based on the surrounding anatomy. These situations bring distinct characteristics and challenges to our dataset. To ensure annotation consistency, we adopt the cross-validation strategy: each clip is initially annotated by four surgeons, followed by a review and refinement process conducted by two additional surgeons. Fig. 2 presents examples of annotations for various bleeding situations. All data has been approved by the ethics statement and will be available.
@@
Data Analysis. 1) Clip Distribution: SurgBlood includes 5,330 frames extracted from 95 video clips. As shown in Fig. 3(a), each clip contains an average of 56 frames, with the longest containing 300 frames and the shortest containing 8 frames. We also counted the distribution of bleeding types: gallbladder (21.64%), cystic triangle (25.01%), vessel (15.78%), and gallbladder bed (37.75%).
@@
2) Bleeding Ratio: We calculate the proportion of frames containing bleeding regions and points. As shown on the left of Fig. 3(b), both bleeding regions and points have a high frame rate, where the slightly higher rate with bleeding regions is due to the partial occlusion of bleeding points.
@@
3) Space Statistics: We analyze the spatial distribution and center bias of bleeding regions and points across all samples in SurgBlood. The right of Fig. 3(c) provides statistical insights into the distances of bleeding region centers and bleeding points from the image center. We can see that bleeding caused by surgical operations predominantly occurs near the image center, and bleeding points are contained within bleeding regions for the vast majority of the time.
@@ 4 Proposed Method
Figure 4: Overview of the proposed BlooDet. Our framework comprises a mask branch and a point branch to jointly detect bleeding regions and bleeding points. Cross-branch guidance and adaptive prompt embedding allow our framework to reach a co-optimized state. Zoom in for details.

### v3

@@ Abstract
Intraoperative bleeding in laparoscopic surgery causes rapid obscuration of the operative field to hinder the surgical process and increases the risk of postoperative complications. Intelligent detection of bleeding areas can quantify the blood loss to assist decision-making, while locating bleeding points helps surgeons quickly identify the source of bleeding and achieve hemostasis in time to improve surgical success rates. To fill the benchmark gap, we first construct a real-world laparoscopic surgical bleeding detection dataset, named SurgBlood, comprising 5,330 frames from 95 surgical video clips with bleeding region and point annotations. Accordingly, we develop a dual-task synergistic online detector called BlooDet, enabling simultaneous detection of bleeding regions and points in laparoscopic surgery. The baseline embraces a dual-branch bidirectional guidance design based on Segment Anything Model 2. The mask branch detects bleeding regions through adaptive edge and point prompt embeddings, while the point branch leverages mask memory to induce bleeding point memory modeling and captures point motion direction via inter-frame optical flow. By coupled bidirectional guidance, our framework explores spatial-temporal correlations while exploiting memory modeling to infer current bleeding status. Extensive experiments indicate that our method outperforms 13 counterparts in bleeding detection. Code and data are available.
@@ 1 Introduction
Minimally invasive surgery has revolutionized clinical healthcare by reducing patient trauma and accelerating postoperative recovery [varghese2024artificial]. However, intraoperative bleeding is an emergency that significantly impacts surgical safety and efficiency in endoscopic surgery [gallaher2022acute]. Rapid changes in the amount and speed of bleeding can severely obscure the surgical field, delaying the surgeon’s response and reducing the success rate of surgery. More prolonged bleeding increases the risk of organ damage and postoperative complications [deziel1993complications]. Therefore, utilizing computer-assisted techniques to detect bleeding regions and localize bleeding points holds significant clinical value. In one respect, detecting bleeding areas can quantify blood loss, providing timely support for intraoperative decision-making. In another respect, precise bleeding point localization enables surgeons to control hemorrhage promptly to ensure safety.
@@
Figure 1: (a): Illustration of bleeding detection task with samples in SurgBlood and predictions of our solution. (b): The proposed BlooDet performs dual-branch bidirectional guidance for synergistic bleeding region and point detection.
@@
Despite the popularity of laparoscopic surgery, automated detection of bleeding regions and bleeding points still faces numerous challenges [sunakawa2024deep]. Due to the narrow field of view under laparoscopy and unstable lighting conditions, the anatomical structures are incompletely exposed, which increases the difficulty of extracting discriminative representations. Additionally, the rapid accumulation and flow of blood can change tissue appearance and infiltrate surrounding tissues, reducing the availability of low-level visual clues and complicating the detection of bleeding regions. The bleeding points may also be buried by blood or obscured by tissues, making it difficult to quickly locate and continuously track [mori2024red]. Beyond these challenges, intelligent bleeding warning involves detecting bleeding regions and locating bleeding points during dynamic surgical procedures [hirai2022evaluation]. This requires a reliable multi-task online detector that models fine-grained spatial-temporal relationships in surgical videos for accurate predictions. Further, the lack of publicly available multi-task real bleeding datasets remains a major obstacle to progress in surgical intelligence.
@@
To advance research on bleeding region and point detection in surgical videos, we construct a new actual laparoscopic surgery bleeding dataset, named SurgBlood. Our dataset comprises a total of 5,330 video frames from 95 laparoscopic video clips, encompassing multiple types and intensities of bleeding during surgery. As displayed in Fig. 1(a), SurgBlood also provides pixel-level annotations of bleeding regions and bleeding point coordinates by hepatobiliary surgeons, supporting the joint detection of bleeding regions and points. We evaluate several task-relevant methods on SurgBlood to establish a comprehensive benchmark for laparoscopic intraoperative bleeding detection, driving further research in surgical assistance.
@@
Existing learning-based methods have been demonstrated to be effective in bleeding region detection [li2020deep, bourbakis2005neural]. However, most algorithms [bourbakis2005neural, li2020deep, das2023multi] are designed for image or keyframe analysis, lacking the ability to model temporal dependencies in surgical videos. In addition, previous methods [su2022spatio, mao2024pitsurgrt, sunakawa2024deep] mainly focus on bleeding region detection, which falls short of addressing the clinical needs in locating the bleeding source. With the emergence of large vision models, Segment Anything Model 2 (SAM 2) [ravi2024sam] unleashed powerful visual representation capabilities for video sequence modeling. Subsequently, a series of SAM 2-based frameworks [liu2024surgical, pei2024evaluation, zhu2024medical, liu2025resurgsam2] have been proposed in the clinical medicine domain, but have not yet been unified into a multi-task paradigm. Multi-task frameworks [das2023multi, mao2024pitsurgrt] can detect both regions and keypoints, but they neglect spatial-temporal modeling, leading to difficulties in stably tracking the movement of bleeding sources. Moreover, current multi-task architectures are unable to fully harness the potential of SAM 2 in cross-task joint optimization.
@@
To meet the clinical demand of bleeding region and point detection, in this paper, we propose a dual-task online baseline model called BlooDet, which embraces a dual-branch bidirectional guidance scheme based on SAM 2 to synergistically optimize both tasks. As illustrated in Fig. 1(b), our framework consists of two branches: Mask branch and Point branch. In the mask branch, we embed an edge generator that performs multi-scale perception of spatial-temporal features with the wavelet Laplacian filter to generate edge prompts, mitigating the problem of blurred bleeding boundaries. Meanwhile, we incorporate bleeding points produced from the point branch as point prompts and combine them with edge prompts to facilitate bleeding region detection. For the point branch, considering the movement of bleeding points within the field of view is influenced by the relative motion of the camera, we leverage inter-frame optical flow and mask memory to estimate camera motion and viewpoint offsets, improving the location accuracy of bleeding points and reducing interference caused by surrounding blood blurring. Further, we integrate mask memory trails from the mask branch to enhance bleeding point perception. By coupling clues and co-guiding across two branches, BlooDet can exploit spatial-temporal associations between bleeding regions and points. Extensive experiments demonstrate that our approach achieves superior performance in both tasks, e.g., 64.88% IoU for bleeding region detection and 83.69% PCK-10% for bleeding point detection. The main contributions are four-fold:
  • We debut intraoperative bleeding region and point detection tasks in surgical videos and contribute a real-world dataset to advance the surgical intelligent assistance.
  • We introduce a dual-task synergistic online detector BlooDet, which adopts a dual-branch structure and performs co-optimization by mutual prompts and bidirectional guidance for bleeding region and point detection.
  • The inter-frame optical flow and mask memory are utilized in point branch for capturing movement cues and providing spatial-temporal modeling. Besides, the edge generator and point prompting strategy in the mask branch exploits multi-scale wavelet Laplacian filters to enhance edge perception, while incorporating bleeding points for mask prompt embedding.
  • We establish a comprehensive benchmark and evaluate 13 task-related models on SurgBlood. Experimental results indicate that our method achieves superior performance for joint bleeding region and point detection.
@@ 2 Related Work
Bleeding Region Detection. Bleeding region detection has been explored across various medical scenarios, such as intracranial hemorrhage detection [li2020deep], capsule endoscopy bleeding recognition [bourbakis2005neural], and retinal hemorrhage identification [wu2024automatic]. Deep learning-based methods employ convolution and attention mechanisms to extract discriminative features for bleeding localization. Sunakawa et al. [sunakawa2024deep] developed a semantic segmentation model for automatically recognizing bleeding regions on the anatomical structure of the liver. Nonetheless, existing methods primarily focus on mask-level bleeding detection, overlooking the localization of the bleeding source. To bridge this gap, we introduce a unified paradigm for the synergistic detection of bleeding regions and points in surgical videos.
@@
Keypoint Detection in Medical Domain. Keypoint detection plays a crucial role in various clinical applications, e.g., pathological site identification and anatomical landmark localization [vorontsov2024foundation, ali2025objective]. Existing keypoint detection methods are usually classified into three categories: 1) Context-aware spatial methods. This technique exploits the stability and uniformity of keypoint spatial distributions to improve localization accuracy [zhang2020headlocnet, liu2019misshapen, tuysuzoglu2018deep, pei2024depth]. 2) Multi-stage learning strategies. These architectures follow a coarse-to-fine process to refine keypoint localization through gradual optimization and integrating shallow and deep layers [zheng20153d, dai2024cephalometric, zhong2019attention]. 3) Multi-task learning frameworks, which jointly optimize image segmentation and keypoint detection by constructing a union network [zhang2020context, das2023multi, mao2024pitsurgrt]. To enhance mutual guidance between tasks, we propose a collaborative dual-branch model that couples bleeding region and point detection.
Figure 2: Illustration of bleeding types in SurgBlood.
Table 1: Summary of publicly available surgical bleeding datasets. SurgBlood provides richer real-world surgical videos with region and point-level annotations, showing holistic bleeding conditions.
Dataset Year Source Type #Videos #Frames Region Point
Rabbani et al. [rabbani2022video] 2022 Human Image 751
HemoSet [miao2024hemoset] 2024 Animal Video 10 857
SurgBlood (Ours) 2025 Human Video 95 5,330
3 SurgBlood Dataset
@@
Currently, there are few publicly available datasets for surgical bleeding detection, as summarized in Table 1. HemoSet [miao2024hemoset] provides bleeding samples based on live animal robotic surgery, including 857 labeled frames for bleeding areas without bleeding points. Rabbani et al. [rabbani2022video] released a dataset of gynecologic laparoscopic surgeries, containing 751 frames of active bleeding rather than video clips. To this end, we construct a brand-new dataset, SurgBlood, specifically for bleeding region and point detection in laparoscopic surgery. We provide an overview of our dataset from the following three aspects: data collection, data annotation, and data analysis.
Figure 3: Statistical distribution of video clips in SurgBlood.
3.1 Data Collection
[…]
Figure 4: Bleeding distribution. Left: proportion of frames with bleeding region and point; Right: Distance of bleeding region center and bleeding point to image center.
Figure 5: Overview of the proposed BlooDet. Our framework comprises a mask branch and a point branch to jointly detect bleeding regions and bleeding points. Cross-branch guidance and adaptive prompt embedding allow our framework to reach a co-optimized state.
3.2 Data Annotation
@@
To ensure the annotation quality of SurgBlood, the invited hepatobiliary surgeons meticulously annotate and review each video clip. During the labeling process, the surgeon uses both static frames and dynamic video sequences to label bleeding regions and bleeding point coordinates for each frame. Annotation is guided by the following principles: 1) For bleeding regions, pooled blood and inactive sparse bloodstains are not annotated. 2) For bleeding points, if the bleeding point is not obscured by tissue or instruments, its coordinates are labeled; if coordinates are difficult to confirm, refer to surrounding anatomy or previous frames to label it. To ensure annotation consistency, we adopt the cross-validation strategy: each clip is initially annotated by four surgeons, followed by a review and refinement process conducted by two additional surgeons. Fig. 2 presents examples of annotations for various bleeding situations. We also provide the analysis of inter-annotator agreement in Supp. H. More statistics and data exhibitions can refer to Supp. B.
3.3 Data Analysis
@@
  • Clip Distribution: SurgBlood includes 5,330 frames extracted from 95 video clips. As shown in Fig. 3, each clip contains an average of 56 frames, with the longest containing 300 frames and the shortest containing 8 frames. We also counted the distribution of bleeding types: gallbladder (21.64%), cystic triangle (25.01%), vessel (15.78%), and gallbladder bed (37.75%).
@@
  • Bleeding Ratio: We calculate the proportion of frames containing bleeding regions and points. As shown on the left of Fig. 4, both bleeding regions and points have a high frame rate, where the slightly higher rate with bleeding regions is due to the partial occlusion of bleeding points.
@@
  • Space Statistics: We analyze spatial distribution of bleeding regions and points across all samples in SurgBlood. The right of Fig. 4 provides statistical insights into the distances of bleeding region centers and bleeding points from the image center. We observed that bleeding primarily occurred in the central zone of surgical manipulation.
@@ 4 Proposed Method
BlooDet is a dual-task collaborative detector for simultaneous bleeding region and point detection in surgical videos. The region detection and point localization tasks are tightly coupled: point localization enforces fine-grained geometric consistency within each predicted region (suppressing loose or false detections), while region detection supplies instance-level spatial context that constrains and stabilizes keypoint regression in cluttered scenes. This mutual dependency implies that the two tasks constrain and optimize each other rather than being solved in isolation. We formalize this interdependence with a coupled objective in which parameters of one branch impact the optimization of the other:
$\{\bm{\theta}^{*},\bm{\vartheta}^{*}\}=\arg\min_{\bm{\theta},\,\bm{\vartheta}}\Big[\mathcal{L}_{\mathtt{m}}\big(\bm{\theta}\!\left(\bm{\vartheta}\right)\big)+\mathcal{L}_{\mathtt{p}}\big(\bm{\vartheta}\!\left(\bm{\theta}\right)\big)\Big],$ (1)
where $\bm{\theta}$ and $\bm{\vartheta}$ denote the model parameters of the region detection and the point localization networks, respectively. The objectives $\mathcal{L}_{\mathtt{m}}$ and $\mathcal{L}_{\mathtt{p}}$ are the training losses associated with these two tasks.